# Multilayer Artificial Benchmark for Community Detection (mABCD)

### v1

@@ Abstract
The Artificial Benchmark for Community Detection (ABCD) model is a random graph model with community structure and power-law distribution for both degrees and community sizes. The model generates graphs similar to the well-known LFR model but it is faster, more interpretable, and can be investigated analytically. In this paper, we use the underlying ingredients of the ABCD model and introduce its variant for multilayer networks, mABCD.
@@ 1 Introduction
One of the most important features of real-world networks is their community structure, as it reveals the internal organization of nodes [24, 35]. In social networks, communities may represent groups by interest; in citation networks, they correspond to related papers; in the Web graph, communities are formed by pages on related topics, etc. Identifying communities in a network is therefore valuable as this information helps us to understand the network structure better.
Unfortunately, there are very few datasets with ground-truth communities identified and labelled. As a result, there is a need for synthetic random graph models with community structure that resemble real-world networks to benchmark and tune clustering algorithms that are unsupervized by nature. The LFR (Lancichinetti, Fortunato, Radicchi) model [42, 41] is a highly popular model that generates networks with communities and, at the same time, allows for heterogeneity in the distributions of both node degrees and of community sizes. It became a standard and extensively used method for generating artificial networks.
@@
A similar synthetic network to LFR, the Artificial Benchmark for Community Detection (ABCD[34] was recently introduced and implemented [^1] https://github.com/bkamins/ABCDGraphGenerator.jl/, including a fast implementation [^2] https://github.com/tolcz/ABCDeGraphGenerator.jl/ that uses multiple threads (ABCDe[38]. Undirected variants of LFR and ABCD produce graphs with comparable properties, but ABCD/ABCDe is faster than LFR and can be easily tuned to allow the user to make a smooth transition between the two extremes: pure (disjoint) communities and random graphs with no community structure. Moreover, it is easier to analyze theoretically — for example, in [33] various theoretical asymptotic properties of the ABCD model are investigated, including the modularity function that, despite some known issues such as the “resolution limit” reported in [25], is an important graph property of networks in the context of community detection. In [5], some interesting and desired self-similar behaviour of the ABCD model is analyzed; namely, it is shown that the degree distribution of ground-truth communities is asymptotically the same as the degree distribution of the whole graph (appropriately normalized based on their sizes). Finally, the building blocks in the model are flexible and may be adjusted to satisfy different needs. Indeed, the original ABCD model was recently adjusted to include potential outliers (ABCD+o[36], overlapping communities (ABCD+o2[4], and extended to hypergraphs (h–ABCD[37] [^3] https://github.com/bkamins/ABCDHypergraphGenerator.jl. The ABCD model is used by practitioners but, for the reasons mentioned above, it also gains recognition among scientists. For example, [2] suggests to use Adjusted Mutual Information (AMI) between the partitions returned by various algorithms with the ground-truth partitions of synthetically generated random graphs, ABCD and LFR. In particular, they use both models to compare 30 community detection algorithms, mentioning that being directly comparable to LFR, ABCD offers additional benefits, including higher scalability and better control for adjusting an analogous mixing parameter. In the context of this paper, the most important of the above features is that the ABCD model is flexible and can be easily extended. In this paper, we utilize that property to deliver a new version producing multilayer networks.
@@
The study of complex networks has evolved significantly over the past two decades, driven by the growing need to model and analyze interconnected systems. Among the notable advancements in this field is the development of multilayer networks, a powerful framework that captures multiple types of relationships within a single network structure. Unlike traditional single-layer networks where all edges represent the same type of interaction, multilayer networks allow nodes to participate in diverse types of relationships across different layers. For instance, in a multilayer social network, layers may represent distinct relations such as friendships, coworker connections, family ties, and online interactions via platforms like Facebook or LinkedIn [40, 21]. This flexibility makes multilayer networks uniquely suited to model the complexity of real-world systems. The concept of multilayer networks was introduced to address the limitations of traditional network models, particularly their inability to represent systems with heterogeneous interactions. Early research focused on formalizing the structure of multilayer networks, including their mathematical representation and classification. Kivelä et al. [40] provided a seminal framework for multilayer networks, categorizing them into multiplex, interconnected, and temporal networks. This foundational work established the groundwork for studying the dynamics and structures of multilayer systems, sparking a surge of interest in the network science community.
@@
As we already mentioned, one of the most important aspects of network analysis is community detection, which aims to identify groups of nodes that are densely connected within the network while being sparsely connected to nodes outside their group. In multilayer networks, this problem is significantly more complex because communities may exist consistently across layers, vary between layers, or even exhibit dependencies across layers [43]. The field of community detection in multilayer networks emerged shortly after the formalization of multilayer network structures, with pioneering studies exploring both theoretical and algorithmic approaches [8, 53, 14]. The survey by Magnani et al. [43] provides an exhaustive overview of existing methods for community detection in multilayer networks, categorizing them into optimization-based methods (e.g., modularity maximization), statistical methods (e.g., stochastic block models), and heuristic approaches. These methods address diverse challenges, such as detecting overlapping communities, identifying temporal changes, and incorporating inter-layer dependencies. Despite these advancements, evaluating community detection algorithms remains challenging due to the lack of real-world multilayer datasets with ground-truth community structures.
To overcome the scarcity of real-world datasets with known community structures, researchers have developed synthetic models capable of generating multilayer networks with predefined community structures. These models provide a controlled environment for testing and validating community detection algorithms, ensuring reproducibility and comparability of results. The models proposed so far, although important milestones have some limitations that can affect their applicability to different research scenarios (see Section 2.1 for details). These limitations underscore the ongoing need for a new model that better replicates the multifaceted nature of real-world multilayer networks and is faster in generating multilayer networks.
@@
In this paper, we use the underlying flexible ingredients of the ABCD family of models and introduce its variant for multilayer networks, mABCD [^4] https://github.com/KrainskiL/MLNABCDGraphGenerator.jl. It is a journal version of a short conceptual paper [13] presented at the 20th Workshop on Modelling and Mining Networks (WAW2025). We significantly extend the previous work by improving, evaluating, and validating the model as well as setting mABCDin the context of related works.
@@
The paper is structured as follows. We start with a gentle introduction to various aspects of multilayer networks (Section 2), including an overview of existing synthetic models, summarizing approaches to measure correlations between layers. Then, we present design assumptions under which mABCD was developed (Section 3). In Section 4, we formally define the mABCD model. Experiments highlighting properties of the model are discussed in Section 5. We finish the paper with a demonstration of how one can use the mABCD model to investigate properties of complex networks and algorithms that are running on them. As an example, we concentrated on the spreading of information (Section 6), which constitutes a problem of greater complexity in the realm of multilayer networks compared to its counterpart in classic graphs.
@@ 2 Related Work
In this section, we introduce the standard notation used across the paper and properties of interest, especially from the perspective of multilayer networks [21, 40]. For a given $n\in{\mathbb{N}}=\{1,2,\ldots\}$, we use $[n]$ to denote the set consisting of the first $n$ natural numbers, that is, $[n]=\{1,2,\ldots,n\}$. We define a multilayer network as a quadruple $M=([n],[\ell],V=[n]\times[\ell],E)$, where
@@
  • $[\ell]$ is a set of layers (for example, different social networking platforms, such as LinkedIn, Facebook and Instagram, on which actors interact with each other),
@@
2.1 Existing Synthetic Models
@@
The first synthetic models for multilayer networks, with known community structure, were extensions of existing single-layer benchmarks. For example, in [11], the mLFR model was introduced, as an extension of the original LFR benchmark, to generate multilayer networks with realistic (scale-free) degree and community size distributions. While the mLFR model preserves realistic degree and community size distributions, it assumes homogeneity in community generation across layers, limiting its applicability to networks where community structures vary between layers. Another drawback is its limited ability to model inter-layer correlations, a critical aspect of many real-world multilayer networks. Furthermore, the mLFR benchmark inherits the computational complexity of the original LFR model, making it resource-intensive for generating large networks with detailed configurations. Another example of early works presenting multilayer network generators is [49], the MuNeG generator was proposed as an extension of the Eldardiry model [22]. It is governed by six scalar parameters: the number of nodes, layers, and groups; the probability that two nodes from the same (or different) group are connected; and community homophily. However, as an early contribution to the field of multilayer network generation, its functionality is limited. The model does not capture inter-layer relations, nor does it provide direct control over the distributions of degrees or communities.
@@
As the body of literature on multilayer networks expanded, various generators addressing specific forms of multilayer networks emerged. For example, [54] introduces a synthetic model designed to construct multilayer networks in which a link between two actors may exist in only one layer, thereby representing a social system composed of four distinct types of relationship: support, sympathy, band, and community. Elsisy et al. [23] also proposed a generator tailored to a specific problem. Specifically, they developed a framework for modelling covert social structures, such as criminal or terrorist networks. Their goal was to provide a model capable to represent both the organizational structure and the community partition of the considered system. These models, despite useful for specific problems, remain limited in terms of their applicability to more general scenarios.
@@
A review of related work also reveals a broad influx of methods extending the Stochastic Block Model [30]. For instance, Huang et al. [31] build upon this framework by assuming that a given community may appear in multiple layers, and propose a method for retrieving such exactly matching groups from the network. On the other hand, the generator proposed by [47] emerges as an indirect outcome of efforts to generalize the modularity function to a multilayer setting. It attempts to address several limitations of extensions of the Stochastic Block Model, such as the restriction to identical communities across layers, but imposes an additional constraint: any inter-layer dependencies are induced solely through the multilayer partition and the coupling of nodes. Another significant contribution (also basing on the Stochastic Block Model) came from Bazzi et al. [6], who proposed a framework for generating networks with mesoscale structures, including communities. This is a very general and powerful model that enables researchers to control the properties of generated networks, such as the strength of inter-layer connections and the overlap of communities between layers. Because of its versatility (multilayer, temporal networks and more), the model might be challenging to use for a novice researcher (the need to define an inter-layer dependency tensor, implementing a new network model if one wants to use something else than the Degree Corrected Stochastic Block Model) and it suffers from high computational complexity (see Figure 15 and Table 4). Finally, it lacks the flexibility needed to reflect the various dependencies existing in real networks (for example, injecting degree sequences from real multilayer networks). As a result, despite its advanced capabilities, the model does not address all needs.
@@
Magnani et al. [43], in the recent review paper on community detection in multiplex networks, have not used the model proposed by Bazzi. Instead, they proposed a simple model for generating multilayer networks with community structures designed specifically for various algorithm comparisons. This model, partially integrated into the multinet library [44], simplifies the generation of multilayer networks and facilitates the benchmarking of various community detection algorithms. While this simplicity facilitates comparisons of different algorithms, it limits the model’s ability to replicate the complexity of real-world networks, such as diversity in community sizes. Furthermore, the model offers only basic representations of inter-layer dependencies, which reduces its relevance for studying networks with varying edge correlations between layers.
@@ 3 Design Assumptions
Before building the model, it is essential to address the question if and how degrees, edges, and partitions correlate between layers in real-world networks. Existing studies [43, 12] have not provided a definitive answer to this question. To bridge this gap, we conducted an analysis of eight real-world networks from diverse domains and of different sizes. In the following section, we first present a methodology of computing these correlations. Then, we discuss obtained results and conclude with the assumptions underlying the design of mABCD.

### v2

@@ Abstract
One of the most persistent challenges in network science is the development of various synthetic graph models to support subsequent analyses. Among the most notable frameworks addressing this issue is the Artificial Benchmark for Community Detection (ABCD) — a random graph model with community structure and power-law distribution for both degrees and community sizes. The model generates graphs similar to the well-known LFR model but it is faster, more interpretable, and can be investigated analytically. In this paper, we use the underlying ingredients of ABCD and introduce its variant, mABCD, thereby addressing the gap in models capable of generating multilayer networks. The uniqueness of the proposed approach lies in its flexibility at both levels of modelling: the internal structure of individual layers and the inter-layer dependencies, which together make the network a coherent structure rather than a collection of loosely coupled graphs. In addition to the conceptual description of the framework, we provide a comprehensive analysis of its efficient Julia implementation. Finally, we illustrate the applicability of mABCD to one of the most prominent problems in the area of complex systems — spreading phenomena analysis.
@@ 1 Introduction
One of the most important features of real-world networks is their community structure, as it reveals the internal organization of nodes [fortunato2010community, kaminski2021mining]. In social networks, communities may represent groups by interest; in citation networks, they correspond to related papers; in the Web graph, communities are formed by pages on related topics, etc. Identifying communities in a network is therefore valuable as this information helps us to understand the network structure better.
The need for synthetic graph models with community structure that resemble real-world networks arises for several reasons. Foremost, when modelling a given system is infeasible (for instance, due to data scarcity), such frameworks can provide a structure that globally approximates the analyzed environment. For example, in the field of epidemiology, the spread of diseases is commonly simulated on synthetic networks owing to the difficulty of obtaining ground-truth data at the human level [watroba2023influence] — only recently have some countries developed network models resembling the social structures of their populations [panayiotou2025swedishstatenetwork]. Continuing this example, simulating disease dissemination using the model proposed by [er-model] could be less accurate than doing so on a network generated by the stochastic block model [holland1983stochastic], as the latter yields community-like structures that better reflect social bonds. Another motivation for developing synthetic graph generators with community structure lies in the need to benchmark and tune clustering algorithms that are unsupervised by nature, given that only a few datasets exist with ground-truth communities identified and labelled [magnani2021community]. Having such models enables experiments to be conducted in a controlled environment, thereby making the assessment of different community detection algorithms more reliable and reproducible. The generators proposed so far, although important milestones, have some limitations that can affect their applicability to different research scenarios (see Section 2.1 for details). These limitations underscore the ongoing need for new models, tailored to both more advanced approaches to modelling complex systems and the increasing size of networks used in analysis.
@@
The LFR (Lancichinetti, Fortunato, Radicchi) model [lancichinetti2008benchmark, lancichinetti2009benchmarks] is a highly popular model that generates networks with communities and, at the same time, allows for heterogeneity in the distributions of both node degrees and of community sizes. It became a standard and extensively used method for generating artificial networks. A similar synthetic network to LFR, the Artificial Benchmark for Community Detection (ABCD[kaminski2021artificial] was recently introduced and implemented [^1] https://github.com/bkamins/ABCDGraphGenerator.jl/, including a fast implementation [^2] https://github.com/tolcz/ABCDeGraphGenerator.jl/ that uses multiple threads (ABCDe[kaminski2022abcde]. Undirected variants of LFR and ABCD produce graphs with comparable properties, but ABCD/ABCDe is faster than LFR and can be easily tuned to allow the user to make a smooth transition between the two extremes: pure (disjoint) communities and random graphs with no community structure. Moreover, it is easier to analyze theoretically — for example, in [kaminski2022modularity] various theoretical asymptotic properties of the ABCD model are investigated, including the modularity function that, despite some known issues such as the “resolution limit” reported in [fortunato2007resolution], is an important graph property of networks in the context of community detection. In [self-similarityABCD], some interesting and desired self-similar behaviour of the ABCD model is analyzed; namely, it is shown that the degree distribution of ground-truth communities is asymptotically the same as the degree distribution of the whole graph (appropriately normalized based on their sizes). Finally, the building blocks in the model are flexible and may be adjusted to satisfy different needs. Indeed, the original ABCD model was recently adjusted to include potential outliers (ABCD+o[kaminski2023artificial], overlapping communities (ABCD+o2[barrett2025artificial], and extended to hypergraphs (h–ABCD[kaminski2023hypergraph] [^3] https://github.com/bkamins/ABCDHypergraphGenerator.jl. The ABCD model is used by practitioners but, for the reasons mentioned above, it also gains recognition among scientists. For example, [aref2022bayan] suggests to use Adjusted Mutual Information (AMI) between the partitions returned by various algorithms with the ground-truth partitions of synthetically generated random graphs, ABCD and LFR. In particular, they use both models to compare 30 community detection algorithms, mentioning that being directly comparable to LFR, ABCD offers additional benefits, including higher scalability and better control for adjusting an analogous mixing parameter. In the context of this paper, the most important of the above features is that the ABCD model is flexible and can be easily extended, as we utilize that property to deliver a new model to producing multilayer networks.
@@
The study of complex networks has evolved significantly over the past two decades, driven by the growing need to model and analyze interconnected systems. Among the notable advancements in this field is the development of multilayer networks, a powerful framework that captures multiple types of relationships within a single network structure. Unlike traditional single-layer (a.k.a. monoplex) networks where all edges represent the same type of interaction, multilayer networks allow nodes to participate in diverse types of relationships across different layers. For instance, in a multilayer social network, layers may represent distinct relations such as friendships, coworker connections, family ties, and online interactions via platforms like Facebook or LinkedIn [kivela2014multilayer, dickison2016multilayer]. This flexibility makes multilayer networks uniquely suited to model the complexity of real-world systems. The concept of multilayer networks was introduced to address the limitations of traditional network models, particularly their inability to represent systems with heterogeneous interactions. Early research focused on formalizing the structure of multilayer networks, including their mathematical representation and classification. [kivela2014multilayer] provided a seminal framework for multilayer networks, categorizing them into multiplex, interconnected, and temporal networks. This foundational work established the groundwork for studying the dynamics and structures of multilayer systems, sparking a surge of interest in the network science community.
@@
As we already mentioned, one of the most important aspects of network analysis is community detection, which aims to identify groups of nodes that are densely connected within the network while being sparsely connected to nodes outside their group [deluca2023communitydetectionapproach]. In multilayer networks, this problem is significantly more complex because communities may exist consistently across layers, vary between layers, or even exhibit dependencies across layers [magnani2021community]. The field of community detection in multilayer networks emerged shortly after the formalization of multilayer network structures, with pioneering studies exploring both theoretical and algorithmic approaches [berlingerio2011finding, tang2012community, brodka2013introduction]. The survey by [magnani2021community] provides an exhaustive overview of existing methods for community detection in multilayer networks, categorizing them into optimization-based methods (e.g., modularity maximization), statistical methods (e.g., stochastic block models), and heuristic approaches. These methods address diverse challenges, such as detecting overlapping communities, identifying temporal changes, and incorporating inter-layer dependencies. Despite these advancements, the problem of community retrieval in multilayer networks remains pertinent, and new approaches continue to emerge [karimi2020multiplexcommunitydetection, qing2024bipartitemixedmembership]. One of the major challenges associated with this task is the evaluation of algorithms, primarily due to the scarcity of real-world multilayer datasets with ground-truth communities. In the aforementioned review study, [magnani2021analysis] recalled only two multilayer networks with partially known partitions, which, however, do not arise from their geometry but from actor labels assigned according to affiliation (e.g., political party). Moreover, both networks were constructed such that their communities are similar (i.e., pillar-like) across layers, which makes them insufficiently representative for benchmarking diverse community detection algorithms.
@@
In this paper, we use the underlying flexible ingredients of the ABCD family of models and introduce its variant for multilayer networks, mABCD [^4] https://github.com/KrainskiL/MLNABCDGraphGenerator.jl [^5] Please note that the name of the proposed method is intended to emphasize its connection to the ABCD model rather than to draw the reader’s attention to benchmarking different community detection methods.. It is a journal version of a short conceptual paper [brodka2025multilayer] presented at the 20th Workshop on Modelling and Mining Networks (WAW2025). We significantly extend the previous work by improving, evaluating, and validating the model as well as setting mABCD in the context of related works. In summary, the main contributions of this work can be expressed in five key points:
  • The proposed model, mABCD, is capable of generating synthetic, scale-free multilayer networks with community structure, where structural properties are shaped both within individual layers and across them.
  • We introduce a latent, biscuit-like layer that supports the process of node assignment into communities and is shared across all layers during the generation process. This structure represents intrinsic actor properties and can be readily replaced with embeddings of real-world entities, thereby enhancing the model’s flexibility.
  • The proposed model extends and strengthens the broader ABCD family of synthetic graph generators, contributing to its coherence and generality.
  • mABCD is an efficient model that outperforms its closest equivalent, multilayerGM, by one order of magnitude in time needed to generate small networks (around 1,000 nodes), with the improvement increasing up to two orders of magnitude for larger networks (around 33,000 nodes).
  • We demonstrate a potential application of the model by addressing the problem of spread controllability in multilayer networks.
@@
The paper is structured as follows. We start with a gentle introduction to various aspects of multilayer networks (Section 2), including an overview of existing synthetic models, summarizing approaches to measure correlations between layers. Then, we present design assumptions under which mABCD was developed (Section 3). In Section 4, we formally define the mABCD model. Experiments highlighting properties of the model are discussed in Section 5. We finish the paper with a demonstration of how one can use the mABCD model to investigate properties of complex networks and algorithms that are running on them. As an example, we concentrate on information spread (Section 6), which constitutes a problem of greater complexity in the realm of multilayer networks compared to its monoplex counterpart.
@@ 2 Related Work
In this section, we introduce the standard notation used across the paper and properties of interest, especially from the perspective of multilayer networks [dickison2016multilayer, kivela2014multilayer]. For a given $n\in{\mathbb{N}}=\{1,2,\ldots\}$, we use $[n]$ to denote the set consisting of the first $n$ natural numbers, that is, $n]=\{1,2,\ldots,n\}$. We define a multilayer network as a quadruple $M=([n],[\ell],V=[n]\times[\ell],E)$, where:
@@
  • $\ell$] is a set of layers (for example, different social networking platforms, such as LinkedIn, Facebook and Instagram, on which actors interact with each other),
@@
2.1 Existing synthetic models for multilayer networks
@@
The first synthetic models for multilayer networks, with known community structure, were extensions of existing single-layer benchmarks. For example, in [brodka2016method], the mLFR model was introduced, as an extension of the original LFR benchmark, to generate multilayer networks with realistic (scale-free) degree and community size distributions. While the mLFR model preserves realistic degree and community size distributions, it assumes homogeneity in community generation across layers, limiting its applicability to networks where community structures vary between layers. Another drawback is its limited ability to model inter-layer correlations, a critical aspect of many real-world multilayer networks. Furthermore, the mLFR benchmark inherits the computational complexity of the original LFR model, making it resource-intensive for generating large networks with detailed configurations. Another example of early works presenting multilayer network generators is [popiel2015muneg], the MuNeG generator that was proposed as an extension of the Eldardiry model [eldardiry2010multi-networkfusion]. It is governed by six scalar parameters: the number of nodes, layers, and groups; the probability that two nodes from the same (or different) group are connected; and community homophily. However, as an early contribution to the field of multilayer network generation, its functionality is limited. The model does not capture inter-layer relations, nor does it provide direct control over the distributions of degrees or communities.
@@
As the body of literature on multilayer networks expanded, various generators addressing specific forms of multilayer networks emerged. For example, [tarvid2017stimulation] introduces a synthetic model designed to construct multilayer networks in which a link between two actors may exist in only one layer, thereby representing a social system composed of four distinct types of relationship: support, sympathy, band, and community. [elsisy2022networkgenerator] also proposed a generator tailored to a specific problem. Specifically, they developed a framework for modelling covert social structures, such as criminal or terrorist networks. Their goal was to provide a model capable to represent both the organizational structure and the community partition of the considered system. These models, despite useful for specific problems, remain limited in terms of their applicability to more general scenarios.
@@
A review of related work also reveals a broad influx of methods extending the Stochastic Block Model [holland1983stochastic]. For instance, [huang2019community] build upon this framework by assuming that a given community may appear in multiple layers, and propose a method for retrieving such exactly matching groups from the network. On the other hand, the generator proposed by [pamfil2019relating] emerges as an indirect outcome of efforts to generalize the modularity function to a multilayer setting. It attempts to address several limitations of extensions of the Stochastic Block Model, such as the restriction to identical communities across layers, but imposes an additional constraint: any inter-layer dependencies are induced solely through the multilayer partition and the coupling of nodes. Another significant contribution (also based on the Stochastic Block Model) came from [bazzi2020framework], who proposed a framework — multilayerGM — for generating networks with mesoscale structures, including communities. This is a very general and powerful model that enables researchers to control the properties of generated networks, such as the strength of inter-layer connections and the overlap of communities between layers. Because of its versatility (multilayer, temporal networks and more), the model might be challenging to use for a novice researcher (the need to define an inter-layer dependency tensor, implementing a new network model if one wants to use something else than the Degree Corrected Stochastic Block Model) and it suffers from high computational complexity (see Figure 17 and Table 4). Finally, it lacks the flexibility needed to reflect the various dependencies existing in real networks (for example, injecting degree sequences from real multilayer networks). As a result, despite its advanced capabilities, the model does not address all needs.
@@
[magnani2021community], in the recent review paper on community detection in multiplex networks, have not used the model proposed by Bazzi. Instead, they proposed a simple model for generating multilayer networks with community structures designed specifically for various algorithm comparisons. This model, partially integrated into the multinet library [magnani2021analysis], simplifies the generation of multilayer networks and facilitates the benchmarking of various community detection algorithms. While this simplicity facilitates comparisons of different algorithms, it limits the model’s ability to replicate the complexity of real-world networks, such as diversity in community sizes. Furthermore, the model offers only basic representations of inter-layer dependencies, which reduces its relevance for studying networks with varying edge correlations between layers.
@@ 3 Design Assumptions
Before building the model, it is essential to address the question if and how degrees, edges, and partitions correlate between layers in real-world networks. Existing studies [magnani2021community, brodka2018quantifying] have not provided a definitive answer to this question. To bridge this gap, we conducted an analysis of eight real-world networks from diverse domains and of different sizes. In the following section, we first present a methodology for computing these correlations. Then, we discuss obtained results and conclude with the assumptions underlying the design of mABCD, which stem from this analysis.